Circular No. 7468

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cy IN MCG +3-40-2 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, re-

port the discovery of an apparent supernova (mag 19.5) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60m automated supernova patrol telescope on July 26.11 UT. SN 2000cy is located at α = 15^{h}32^{m}14^{s}.75, δ = +18^{o}40′22′′.9 (equinox 2000.0), which is 29′′.4 west and 5′′.7 south of the center of MCG +3-40-2. The new object was also present on an unfiltered CCD frame taken on July 28.09, but does not appear on Palomar Sky Survey images taken on 1992 June 29 (limiting mag ∼ 21.3) and on 1986 Apr. 12 (limiting mag ∼ 20.2).

COMET C/1999 S4 (LINEAR) J. Licandro, G. Tessicini, and I. Perez, Centro Galileo Galilei; and S.

Hidalgo, Instituto de Astrofísica de Canarias, report that the inner coma of C/1999 S4 appears extremely elongated in the tailward direction on J, H, and K_{s} images obtained with the 3.6-m Telescopio Nazionale Galileo (+ ARNICA camera) at La Palma on July 26.9 and 27.9 UT. There is no clear central condensation, but the photometric peak appears to move from the anti-tailward border of the coma toward p.a. 81^{o} on July 26.9 and 84^{o} on July 27.9, with a velocity of 7′′.4/day (26 m/s). The brightness of the central coma decreased from J = 8.83 ± 0.04 on July 26.9 to 9.46 ± 0.07 on July 27.9, as measured in a 30′′-diameter aperture around the peak. Color maps do not show any major structure, the mean colors being J − K = +0.61 ± 0.05, and H −K = +0.12 ± 0.06 on July 26.9. These data suggest that a major event has occurred in the nucleus of the comet.

A. V. Filippenko and R. Chornock, University of California at Berkeley, report that inspection of a CCD spectrum (range 320–1000 nm) of this comet obtained on July 28 UT with the Shane 3-m reflector at Lick Observatory reveals no clear emission lines superposed on the solar reflection spectrum, unlike the case in spectra they obtained with the same instrument on July 6 and June 27. There is no evidence for the normally strong CN emission near 380 nm. The comet’s morphology is peculiar, lacking an obvious head that is brighter than the adjacent tail, although the leading edge of the head/tail combination is very sharply defined.

S. Nakano, Sumoto, Japan, writes that A. Asami (Bisei Spaceguard Center) was unable to find a peak in the comet’s central-condensation brightness to determine an astrometric position from a CCD image taken in fair conditions on July 28.490 UT with a 0.25-m f/5 reflector.

2000 July 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green